250g of LMnO4 reacts with excess Fe and Hal according two the reaction:

3KMnO4 + 5Fe + 24HCl --> 5FeCl3 + 3MnCl2 + 3KCl + 12H2O

The theoretical yield of FeCl3 is (show your work)

a. 16g

b. 2.6g

c. 260g

d. 430g

e. 1300g
Answer #1

mass of KMnO4 used = 250 gram , molar mass of KMnO4 = 158.034 gram/mole

no of moles = mass in gram/molar mass , substitute values to find no of moles of KMnO4 used

no of moles of KMnO4 used = 250/158.034 = 1.582 moles

balanced chemical reaction is given bellow

3KMnO4 + 5Fe + 24HCl   5FeCl3 + 3MnCl2 + 3KCl

according to reaction 3 moles of KMnO4 produces 5 moles of FeCl3  

therefore 1.582 mole of KMnO4 will produce = (1.582 x 5)/3 = 2.636 moles of FeCl3

no of moles of FeCl3 produced = 2.636 moles , molar mass of FeCl3 = 162.2 gram/mole

mass in gram = no of moles x molar mass , substitute values to find grams of FeCl3 produced

mass of FeCl3 produced in gram = 2.636 x 162.2 = 427.56 gram = theoretical yield

above calculated value of theoretical yield is nearest value 430 which is value of option d therefore option d is correct option

ans d . 430 gram